OpenStudy (littlet):

Could someone help me out? Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 4 inches, and the height of the cone is 6 inches. If the diameter of the bubble gum is 0.8 inches, which of the following can be used to calculate the volume of the cone that can be filled with flavored ice? 1 over 3(3.14)(6^2)(4) − 4 over 3(3.14)(0.43^3) 1 over 3(3.14)(4^2)(6) − 4 over 3(3.14)(0.43^3) 1 over 3(3.14)(6^2)(4) − 4 over 3(3.14)(0.83^3) 1 over 3(3.14)(4^2)(6) − 4 over 3(3.14)(0.8

OpenStudy (littlet):

The last answer cut off (0.83^3)

OpenStudy (littlet):

@clanger555 could you help?

OpenStudy (littlet):

@Aveline

OpenStudy (alekos):

that would be the volume of the cone minus the volume of the bubble gum

OpenStudy (littlet):

How would i find the volume

Arcadiouse (arcadiouse):

V=πr2h 3

OpenStudy (alekos):

volume of a cone is \[(1/3) \pi r ^{2}h\]

OpenStudy (littlet):

1/3 =3.14 4^2 6 is that right?

OpenStudy (alekos):

Yes that's the volume of the cone where r=4 and h=6

OpenStudy (alekos):

Volume of a sphere is ... \[(4/3)\pi r ^{3}\]

OpenStudy (littlet):

for the bubble gum it is only giving diameter

Arcadiouse (arcadiouse):

Radius is half of Diameter

OpenStudy (alekos):

well the diameter is just twice the radius

OpenStudy (littlet):

oh yeah.. i forgot lol

OpenStudy (alekos):

So now you should be able to pick the correct answer

OpenStudy (littlet):

so 4/3 (3.14) .4^3

OpenStudy (littlet):

so B? @alekos

OpenStudy (littlet):

@alekos is B right?

OpenStudy (alekos):

Yes. but I don't know where the 0.43 came from , probably a typo

OpenStudy (littlet):

Yeah it is sorry. Thank you so much!

OpenStudy (alekos):

no worries